### v2

@@ 2.5 Determination of the Number of Factors
In this section, we only introduce two commonly used ways, one is based on the well-known information criterion (IC) method in Bai and Ng (2002), which estimates the number of factors $r$ by
$\widehat{r}=\arg\min_{1\leq k\leq K}\log\left(\frac{1}{NT}\|{\mathbf{X}}-\widehat{\mathbf{F}}_{k}\widehat{\boldsymbol{\Lambda}}_{k}^{\prime}\|_{F}^{2}\right)+k\frac{N+T}{NT}\log(\frac{NT}{N+T}),$ (12)
where $K$ is a prescribed upper bound, $\widehat{\mathbf{F}}_{k}$ and $\widehat{\boldsymbol{\Lambda}}_{K}$ are the estimated factors and loading matrix based on $k$ factors. The other method is based on the eigenvalue ratios introduced by Lam and Yao (2012) and Ahn and Horenstein (2013). Specifically, let $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq...\geq\widehat{\lambda}_{T}$ be the $T$ sample eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$, we adopt the ratio-based method to estimate $r$ by
@@
2.6 Choice of the Tuning Parameters
@@
In practice, the cardinality of each column in ${\mathbf{F}}$ is unknown, and we may choose them by the cross-validation method which is commonly used in machine learning literature. In fact, even though the sparsity parameters $s_{1},...,s_{r}$ of $r$ factor processes may not be the same, it is still a convenient way to assume that $s_{1}=...=s_{r}=s_{0}\asymp T^{\alpha}$ which can simplify the cross-validation procedure significantly. According to the theoretical results in Section 3 and the proofs in the Appendix, when the $r$ sparsity parameters are distinct, the convergence of the estimator for $s^{*}$ can be achieved, where $s^{*}=\max(s_{1},...,s_{r})$. Therefore, we assume that the sparsity parameters of each column are the same in this section. The cross-validation procedure is as follows.
For each fixed $s\in\mathbb{S}$ where $\mathbb{S}$ is a chosen candidate set for the sparsity parameter $s$, we first randomly divide the panel into two segments consisting of one training sample with $N_{1}$ components and the other a testing one with $N_{2}$ components, where $N_{1}\asymp N_{2}\asymp N$ and $N_{1}+N_{2}=N$. Equivalently, we partition the data matrix $\widetilde{\mathbf{X}}\in R^{T\times N}$ into $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\mathbf{X}}_{2}$, where $\widetilde{\mathbf{X}}_{1}\in R^{T\times N_{1}}$ and $\widetilde{\mathbf{X}}_{2}\in R^{T\times N_{2}}$. Let $\widetilde{\mathbf{F}}_{1}^{s}$ be the estimated factors based on the training sample $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\boldsymbol{\Lambda}}_{1}$ be the estimated loading matrix. Define the testing error as
@@
$PC(s)=R(s,\widetilde{\mathbf{F}}_{1}^{s})+r\frac{s}{T^{\alpha}}g(N_{1},T),$
Then, $\widehat{s}$ is the one that minimizes the above criterion, that is,
$\widehat{s}=\arg\min_{s\in\mathbb{S}}PC(s)=\arg\min_{s\in\mathbb{S}}\{R(s,\widetilde{\mathbf{F}}_{1}^{s})+r\frac{s}{T^{\alpha}}g(N_{1},T)\},$ (15)
implying that we choose the optimal $s$ which results in minimal testing errors in the cross-validations. Note that the sparsity strength parameter $\alpha$is unknown and must be specified prior to applying the criterion. Its value depends on the specific application and involves an element of subjective judgment. Importantly, $\alpha$ does not influence the study of risk factors over time, as it only reflects the relative strengths of the nonzero factors. In practice, we recommend setting $\alpha=1/2$, which is sufficient for identifying a robust set of systematic risk events and factors in empirical analyses.
There are many choices for the penalty function $g(N_{1},T)$ as shown in Theorem 4 in Section 3 below. We will discuss the required properties of $g(N_{1},T)$ such that we can establish the consistency of $\widehat{s}$ in theory.
In practice, we may replace the candidate set $\mathbb{S}$ by $smin\leq s\leq smax$ where $smin$ and $smax$ are some prescribed integers. Note that the cross-validation method is different from the classical one in the machine learning literature where the samples are partitioned along the time horizon, whereas we partition the samples over the cross-section in this paper. In fact, we can even perform more cross-validation experiments in order to obtain an optimal parameter $s$. For example, we can choose a sufficiently large integer $J>0$, and we can obtain $J$ testing errors as that in (14) for each fixed $s$ by partitioning the samples for $J$ times. We may denote the testing error by $R_{j}(s,\widetilde{\mathbf{F}}_{1,j}^{s})$ for the $j$-th random partition and the average testing error for $s$ is defined as
@@ 3 Theoretical Properties
In this section, we present the consistency and asymptotic bounds of the estimators from Section 2 under the framework where $N,T\rightarrow\infty$. Assuming the number of factors $r$ is fixed, and $s_{i}$ can be either fixed or slowly increasing, we first derive the asymptotic results under known $r$ and $s_{i}$. We also demonstrate the consistency of the estimators for $r$ and $s_{i}$ thereafter. The mathematical proofs are provided in the Appendix.
@@
There exists a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$ with $s_{i}\asymp T^{\alpha_{i}}$ such that $\|{\mathbf{F}}\|_{0}\leq{\mathbf{s}}$, where ${\mathbf{F}}$ is defined as that in (3) and $\alpha_{i}\in(0,1]$ for $1\leq i\leq r$.
@@
The eigenvalues of $\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}$ are distinct and bounded away from 00 and $\infty$ as $N\rightarrow\infty$.
@@
$e_{i,t}=\sigma_{i}{{\varepsilon}}_{i,t}$ for some $0<c_{1}\leq\sigma_{i}\leq c_{2}<\infty$, where ${{\varepsilon}}_{i,t}$ is independent and identically distributed over $i$ and $t$.
@@
$\underline{{\mathbf{f}}}_{i}$ and $\underline{{\mathbf{e}}}_{j}$ are sub-Gaussian variables for $i=1,...,r$ and $j=1,...,N$ in the sense that
$P(|{\mathbf{v}}^{\prime}\underline{{\mathbf{f}}_{i}}|>x)\leq c\exp(-cx^{2})\quad\text{and}\quad P(|{\mathbf{v}}^{\prime}\underline{{\mathbf{e}}_{i}}|>x)\leq c\exp(-cx^{2}),$
@@
Assumption 1 is standard to characterize the dynamic dependence of the factor processes. See, for example, the Lemma 1 in the appendix of Gao et al. (2019). Assumptions 2-3 establish an identification condition and a sparsity structure for the true factor process ${\mathbf{F}}$. While the factors are sparse, Assumption 2 is not restrictive, as it is reasonable to assume $s_{i}$ diverges and Assumption 2 naturally holds under Assumption 3 by setting $\alpha_{i}=1$ and specifying $s_{i}/T=c_{i}\in(0,1)$, which is in line with the setting in random matrix theory. Note that we could adopt the framework in Uematsu and Yamagata (2022), which treats factors as weak with varying strengths. However, specifying these strengths in a given application can be challenging. Instead, we follow the strong factor framework in Kristensen (2017) under sparsity, and the strength of nonzero factors may diverge if $\alpha_{i}<1$ in Assumption 3. This divergence is merely a matter of the data’s magnitude and does not affect our goal of identifying nonzero factors over time. Assumption 4 specifies that the top eigenvalues of the covariance of the panel data are distinct in order to avoid multiplicity of eigenvalues. Assumptions 2 and 4 provide identifiability for the Model (3) so that the factors and the loading matrix can be uniquely determined asymptotically. See Bai and Ng (2013) for a detailed argument.Assumption 5 is primarily introduced to simplify the theoretical derivations. It can be relaxed to weaker conditions, such as Conditions (c.i)–(c.iii) in Bai and Ng (2013), or replaced with strongly mixing conditions in both spatial and temporal dimensions. Similar independence assumptions have also been adopted in Huang et al. (2022) to streamline theoretical analyses. Assumption 7 is used to derive the Bernstein-type concentration inequality in establishing the convergence rates of the estimated factors, and it can be relaxed to much weaker ones as those in Merlevède et al. (2011).
@@
Let Assumptions 17 hold and $\widehat{\mathbf{v}}_{1}$ be the first solution of (6). As $N,T\rightarrow\infty$, it holds that
@@
To maintain clear interpretation, we continue to use $s_{1}$ instead of $s_{1}\asymp T^{\alpha_{1}}$ in the theorem. From Theorem 1, we see that the angle between the estimated direction $\widehat{\mathbf{v}}_{1}$ and the true one ${\mathbf{v}}_{1}$ is asymptotically equal to zero or $\pi$ if $s_{1}\log(T)/(NT)\asymp\log(T)/(NT^{1-\alpha_{1}})\rightarrow 0$. In fact, we can rewrite $\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}$ as $\cos(\theta)$ where $\theta$ is the angle between $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$. Some remarks for Theorem 1 are as follows.
@@
$\|\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}\|_{2}=\sqrt{2(1-\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})}\leq\sqrt{2(1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2})}=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{NT}}+\frac{s_{1}\log(T)}{NT}+\frac{1}{T}\right),$
and therefore,
$\|\underline{\widehat{\mathbf{f}}}_{1}-\underline{{\mathbf{f}}}_{1}\|_{2}=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{N}}+\frac{s_{1}\log(T)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
implying that we can uniquely determine the factors asymptotically if $s_{1}\log(T)=o(N)$ as $N,T\rightarrow\infty$.
(iii) When the sparsity is proportional to the sample size, i.e., $s_{1}\asymp T$, we have that
@@
Note that $D(\mathcal{M}({\bf H}_{1}),\mathcal{M}({\bf H}_{2}))\in[0,1].$ It is equal to 00 if and only if $\mathcal{M}({\bf H}_{1})=\mathcal{M}({\bf H}_{2})$, and to $1$ if and only if $\mathcal{M}({\bf H}_{1})\perp\mathcal{M}({\bf H}_{2})$. By Lemma A1(i) in Pan and Yao (2008), $D(\cdot,\cdot)$ is a well-defined distance measure on some quotient space of matrices. Alternatively, we may also adopt the measure
@@
Let Assumptions 17 hold and $\widehat{\mathbf{V}}$ be the matrix consisting of the estimated sparse eigenvectors obtained by Algorithm 2 . As $N,T\rightarrow\infty$, it holds that
@@
implying that all the principal angles will be asymptotically zero or $\pi$ if $s^{*}\log(T)=o(NT)$. On the other hand, if all directions of the estimated eigenvectors are in the same direction of the corresponding true ones.
@@
where we used the inequality $(1-\sigma_{i})\leq(1-\sigma_{i})(1+\sigma_{i})=(1-\sigma_{i}^{2})$. Therefore, by (3), we have that
$\|\widehat{\mathbf{V}}-{\mathbf{V}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(T)}{NT}}+\frac{s^{*}\log(T)}{NT}+\frac{1}{T}\right).$
Consequently, we obtain the result for $\widehat{\mathbf{F}}$ as
$\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\sqrt{T}\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(T)}{N}}+\frac{s^{*}\log(T)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
@@
(i) There exits a rotation matrix ${\mathbf{H}}_{s}$ such that
$\max_{1\leq i\leq N}\|\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}\|_{2}=O_{p}(\sqrt{\frac{s^{*}{{}^{2}}\log(N)}{NT^{2}}}\log(T)+\sqrt{\frac{\log(N)}{T}}),$ (22)
where ${\mathbf{H}}_{s}=(\widehat{\mathbf{F}};\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}$ and $s^{*}=\max(\alpha_{1},...,\alpha_{r})$. Furthermore, if $\sqrt{\frac{1}{NT^{1-2\alpha}}}\log(T)=o(1)$ where $\alpha=\max(\alpha_{1},...,\alpha_{r})$, then
@@
(ii) For $1\leq i\leq N$, if $\sqrt{\frac{1}{NT^{1-2\alpha}}}\log(T)=o(1)$, then there exists a rotation matrix ${\mathbf{H}}_{s}$ such that
@@
where ${\mathbf{Q}}$ is the limit of $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T$, and $\boldsymbol{\Gamma}_{i}=\lim_{T\rightarrow\infty}\textnormal{Var}(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}{\mathbf{f}}_{t}e_{i,t})$.
@@
Furthermore, we discuss the consistency of the estimated number of factors. Given the convergence results in Theorems 1 and 2, it is not hard to see that the estimated number of factors are consistent no matter whether we apply the information criterion of Bai and Ng (2002) or the eigenvalue ratio based method in Lam and Yao (2012) and Ahn and Horenstein (2013). The detailed proofs are the same as the techniques used therein, and we omit them to save space here.
Finally, we provide the consistency of the cross-validation method in estimating the sparsity parameters $s_{1},...,s_{r}$. We consider the case where the sparsity parameters of each column of ${\mathbf{F}}$ are the same, i.e., $s_{1}=...=s_{r}=s_{0}$, as discussed in Section 2.6. Otherwise, the following theorem states the consistency for estimating the largest sparsity parameter among the $r$ columns.
@@
Let Assumptions 17 in the Appendix hold. Suppose $\widehat{s}$ is the solution in (15). If $g(N_{1},T)\rightarrow 0$ and $C_{N_{1}T}^{-2}g(N_{1},T)\rightarrow\infty$ as $N,T\rightarrow\infty$, we have
$\lim_{N,T\rightarrow\infty}P(\widehat{s}=s_{0})=1,$
@@
(i) When the columns of ${\mathbf{F}}$ have distinct sparsity parameters, it is an important step to estimate the largest one among $\{s_{1},...,s_{r}\}$ first, and $\widehat{s}$ in Theorem 4 is an estimator for $s^{*}=\max\{s_{1},...,s_{r}\}$. In other words, we can show that
$\lim_{N,T\rightarrow\infty}P(\widehat{s}=s^{*})=1.$ (24)
@@
(ii) In practice, there are many choices for $g(N_{1},T)$ and $\alpha$ in (15). For example, we may take
$g(N_{1},T)=\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T}),$
and $\alpha=1/2$, and
$IC(s)=\ln(R^{J}(s))+r\frac{s}{\sqrt{T}}\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T})$ (25)
can consistently estimate $s_{0}$, where $r$ can be replaced by $\widehat{r}$ obtained in (13).

### v3

@@ 2.3 Determination of the Number of Factors
We introduce two commonly used methods. The first is an information criterion (IC) approach adapted from Bai and Ng (2002), with an added $\log(T)$ penalty to account for the scale of $s$. The number of factors $r$ is estimated by
$\widehat{r}=\arg\min_{1\leq k\leq K}\log\left(\frac{1}{NT}\|{\mathbf{X}}-\widehat{\mathbf{F}}_{k}\widehat{\boldsymbol{\Lambda}}_{k}^{\prime}\|_{F}^{2}\right)+k\frac{N+T}{NT}\log(T)\log\left(\frac{NT}{N+T}\right),$ (11)
where $K$ is a pre-specified upper bound, and $\widehat{\mathbf{F}}_{k}$, $\widehat{\boldsymbol{\Lambda}}_{k}$ are the estimated factors and loadings for $k$ factors. The other method is based on the eigenvalue ratios introduced by Lam and Yao (2012) and Ahn and Horenstein (2013). Specifically, let $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq...\geq\widehat{\lambda}_{T}$ be the $T$ sample eigenvalues of ${\mathbf{X}}{\mathbf{X}}^{\prime}$, we adopt the ratio-based method to estimate $r$ by
@@
2.4 Cross-Sectional Cross-Validation
@@
For theoretical purposes, we may assume $s_{i}/T=\delta_{i}\in(0,1)$, analogous to the dimensional asymptotics commonly considered in random matrix theory. In practice, the cardinality of each column in $\mathbf{F}$} is unknown, and we may choose them by the cross-validation method which is commonly used in machine learning literature. In fact, even though the sparsity parameters $s_{1},...,s_{r}$ of $r$ factor processes may not be the same, it is still a convenient way to assume that $s_{1}=...=s_{r}=s_{0}\asymp T$ which can simplify the cross-validation procedure significantly. According to the theoretical results in Section 3 and the proofs in the Supplement, when the $r$ sparsity parameters are distinct, the convergence of the estimator for $s^{*}$ can be achieved, where $s^{*}=\max(s_{1},...,s_{r})$. Once $s^{*}$ is estimated in this scenario, it can be fixed, allowing the estimation of the second-largest sparsity parameter using the proposed method. For simplicity, in this section, we assume that the sparsity parameters for each column are identical. The cross-validation procedure is outlined as follows.
For each fixed $s\in\mathbb{S}$ where $\mathbb{S}$ is a chosen candidate set for the sparsity parameter $s$, we first randomly divide the panel into two segments consisting of one training sample with $N_{1}$ components and the other a testing one with $N_{2}$ components, where $N_{1}\asymp N_{2}\asymp N$ and $N_{1}+N_{2}=N$. Equivalently, we partition the data matrix ${\mathbf{X}}\in R^{T\times N}$ into $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\mathbf{X}}_{2}$, where $\widetilde{\mathbf{X}}_{1}\in R^{T\times N_{1}}$ and $\widetilde{\mathbf{X}}_{2}\in R^{T\times N_{2}}$. Let $\widetilde{\mathbf{F}}_{1}^{s}$ be the estimated factors based on the training sample $\widetilde{\mathbf{X}}_{1}$ and $\widetilde{\boldsymbol{\Lambda}}_{1}$ be the estimated loading matrix. Define the testing error as
@@
$PC(s)=R(s,\widetilde{\mathbf{F}}_{1}^{s})+rC_{T}\frac{s}{T}g(N_{1},T),$
for some $C_{T}>0$ which will be specified later. Then, $\widehat{s}$ is defined as
$\widehat{s}=\arg\min_{s\in\mathbb{S}}PC(s)=\arg\min_{s\in\mathbb{S}}\{R(s,\widetilde{\mathbf{F}}_{1}^{s})+rC_{T}\frac{s}{T}g(N_{1},T)\},$ (14)
implying that we choose the optimal $s$ which results in minimal testing errors in the cross-validations. There are many choices for the penalty function $g(N_{1},T)$ as shown in Theorem 4 of Section 3 below. We will discuss the required properties of $g(N_{1},T)$ such that we can establish the consistency of $\widehat{s}$ in theory.
In practice, we may restrict the candidate set $\mathbb{S}$ to the range $c_{1}T\leq s\leq c_{2}T$, where $c_{1},c_{2}\in(0,1)$ are prescribed constants. We may choose $C_{T}=\log(T)$ in order to balance the magnitude of the scale of $s$. Note that the cross-validation method is different from the classical one in the machine learning literature where the samples are partitioned along the time horizon, whereas we partition the samples over the cross-section in this paper. In fact, we can even perform more cross-validation experiments in order to obtain an optimal parameter $s$. For example, we can choose a sufficiently large integer $J>0$ and obtain $J$ testing errors as that in (13) for each fixed $s$ by partitioning the samples for $J$ times. We may denote the testing error by $R_{j}(s,\widetilde{\mathbf{F}}_{1,j}^{s})$ for the $j$-th random partition and the average testing error for $s$ is defined as
@@ 3 Theoretical Properties
In this section, we present the consistency and asymptotic bounds of the estimators from Section 2 as $N,T\rightarrow\infty$. We first derive the asymptotic results assuming known $r$ and $s_{i}$, and then establish the consistency of their estimators. Mathematical proofs are provided in the Supplement.
@@
There exists a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$ with $s_{i}/T=\delta_{i}\in(0,1)$ for $1\leq i\leq r$ such that $\|{\mathbf{F}}\|_{0}\leq{\mathbf{s}}$, where $\mathbf{F}$} is defined as that in (3).
@@
The eigenvalues of $\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}$ are distinct and bounded away from 0 and $\infty$ as $N\rightarrow\infty$.
@@
$e_{i,t}=\sigma_{i}{{\varepsilon}}_{i,t}$ for some $0<\underline{\sigma}\leq\sigma_{i}\leq\bar{\sigma}<\infty$, where ${{\varepsilon}}_{i,t}$ is independent and identically distributed over $i$ and $t$.
@@
$\underline{{\mathbf{f}}}_{i}$ and $\underline{{\mathbf{e}}}_{j}$ are sub-Gaussian variables for i=1,,ri=1,...,r and j=1,,Nj=1,...,N in the sense that
$P(|{\mathbf{v}}^{\prime}\underline{{\mathbf{f}}_{i}}|>x)\leq c\exp(-cx^{2})\quad\text{and}\quad P(|{\mathbf{v}}^{\prime}\underline{{\mathbf{e}}_{j}}|>x)\leq c\exp(-cx^{2}),$
@@
Assumption 1 is standard to characterize the dynamic dependence of the factor processes. See, for example, the Lemma 1 in the appendix of Gao et al. (2019). Assumptions 2-3 establish an identification condition and a sparsity structure for the true factor process $\mathbf{F}$}. While the factors are sparse, Assumption 2 is not restrictive, as it is reasonable to assume that $s_{i}$ diverges. Moreover, the factors can be normalized such that Assumption 2 holds under the condition $s_{i}/T=\delta_{i}\in(0,1)$, as specified in Assumption 3, which resembles the setting commonly adopted in random matrix theory such as those in Bai and Silverstein (2010). Alternatively, one could adopt the framework of Uematsu and Yamagata (2022), which models factors as weak with varying strengths. However, specifying the strength levels in practice can be challenging. Instead, we adopt the random matrix theory perspective and assume that the sparsity parameter grows at a fractional rate relative to the time dimension $T$. This is purely a theoretical device and does not impact our primary goal of identifying nonzero factors over time. Assumption 4 specifies that the top eigenvalues of the covariance of the panel data are distinct in order to avoid multiplicity of eigenvalues. Assumptions 2 and 4 provide identifiability for the Model (3) so that the factors and the loading matrix can be uniquely determined asymptotically. See Bai and Ng (2013) for a detailed argument. Assumption 5 is primarily imposed to simplify theoretical derivations. It can be relaxed to weaker conditions, such as Conditions (c.i)–(c.iii) in Bai and Ng (2013), or replaced with strong mixing assumptions in both spatial and temporal dimensions, allowing the use of the Bernstein-type inequality in Merlevède et al. (2011) for the analysis. Similar independence assumptions are also found in Onatski (2012) and Assumption 3 of Huang et al. (2022). Assumption 7 facilitates the derivation of a Bernstein-type concentration bound for establishing convergence rates of the estimated factors, and can likewise be weakened following Merlevède et al. (2011).
@@
Let Assumptions 17 hold and $\widehat{\mathbf{v}}_{1}$ be the first solution of (5). As NN and $T$ increase, it holds that
@@
To maintain clear interpretation, we continue to use $s_{1}$ instead of $s_{1}\asymp T$ in the theorem. From Theorem 1, we see that the angle between the estimated direction $\widehat{\mathbf{v}}_{1}$ and the true one ${\mathbf{v}}_{1}$ is asymptotically equal to zero or $\pi$ if $s_{1}\log(T)/(NT)\asymp\log(T)/N\rightarrow 0$. In fact, we can rewrite $\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}$ as $\cos(\theta)$ where $\theta$ is the angle between $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$. Some remarks for Theorem 1 are as follows.
@@
$\|\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}\|_{2}=\sqrt{2(1-\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})}\leq\sqrt{2(1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2})}=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{NT}}+\frac{s_{1}\log(T)}{NT}+\frac{1}{T}\right).$
Consequently,
@@
Note that $D(\mathcal{M}({\bf H}_{1}),\mathcal{M}({\bf H}_{2}))\in[0,1].$ It is equal to 0 if and only if $\mathcal{M}({\bf H}_{1})=\mathcal{M}({\bf H}_{2})$, and to $1$ if and only if $\mathcal{M}({\bf H}_{1})\perp\mathcal{M}({\bf H}_{2})$. By Lemma A1(i) in Pan and Yao (2008), $D(\cdot,\cdot)$ is a well-defined distance measure on some quotient space of matrices. Alternatively, we may also adopt the measure
@@
Let Assumptions 17 hold and $\widehat{\mathbf{V}}$ be the matrix consisting of the estimated sparse eigenvectors obtained by Algorithm 2. As NN and $T$ increase, it holds that
@@
implying that all the principal angles will be asymptotically zero or $\pi$ if $s^{*}\log(T)=o(NT)$.
@@
where we used the inequality $(1-\sigma_{i})\leq(1-\sigma_{i})(1+\sigma_{i})=(1-\sigma_{i}^{2})$. Consequently, by (3) and Assumption 3, we have that
$\frac{1}{\sqrt{T}}\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{\log(T)}{N}}+\frac{1}{T}\right),$
@@
(i) There exists a rotation matrix ${\mathbf{H}}_{s}$ such that
$\max_{1\leq i\leq N}\|\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}\|_{2}=O_{p}(\frac{\sqrt{\log N}}{N}+\sqrt{\frac{\log N}{T}}),$ (21)
where ${\mathbf{H}}_{s}=(\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}$. Furthermore, if ${\sqrt{T}}/{N}=o(1)$, then
@@
(ii) For $1\leq i\leq N$, if ${\sqrt{T}}/{N}=o(1)$, then there exists a rotation matrix ${\mathbf{H}}_{s}$ as above such that
@@
where $\mathbf{Q}$} is the limit of $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T$, and $\boldsymbol{\Gamma}_{i}=\lim_{T\rightarrow\infty}\textnormal{Var}(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}{\mathbf{f}}_{t}e_{i,t})$ is the long-run covariance matrix.
@@
The following corollary establishes the consistency of the estimated number of factors obtained via (11) or (12).
Corollary 1.
Let Assumptions 17 hold. Suppose $N\asymp T$, $\|{\mathbf{e}}{\mathbf{e}}^{\prime}/N\|_{2}\leq\bar{c}<\infty$, and the $K$-th largest eigenvalue satisfies $\|{\mathbf{e}}{\mathbf{e}}^{\prime}/N\|_{K}\geq\underline{c}>0$, where $K=\min(N,T)/3$ as in (11)–(12). Then both the information criterion in (11) and the ratio-based method in (12) consistently estimate $r$.
Finally, we provide the consistency of the cross-validation method in estimating the sparsity parameters $s_{1},...,s_{r}$. We consider the case where the sparsity parameters of each column of $\mathbf{F}$} are the same, i.e., $s_{1}=...=s_{r}=s_{0}$, as discussed in Section 2.4. Otherwise, the following theorem states the consistency for estimating the largest sparsity parameter among the $r$ columns. Since $s_{0}\asymp T$, it is generally challenging to derive consistency of $\widehat{s}$ when both $N$ and $T$ grow. Therefore, we restrict attention to the asymptotic regime where $N\to\infty$ and $T$ is large but fixed.
@@
Let Assumptions 17 hold. Suppose $s_{0}=c_{0}T$ and $\widehat{s}$ is the solution in (14) with $\mathbb{S}=[c_{1}T,c_{2}T]$ where $0<c_{1}\leq c_{0}\leq c_{2}<1$. If $C_{T}g(N_{1},T)\rightarrow 0$ and $C_{N_{1}T}^{-2}C_{T}g(N_{1},T)\rightarrow\infty$ as $N\rightarrow\infty$ with large $T$, where $C_{NT}=\sqrt{{\log(T)}/{N}}+{1}/{T}$, we have
$\lim_{N\rightarrow\infty}P(\widehat{s}=s_{0})=1,$
@@
(i) Theorem 4 shows consistency of $\widehat{s}$ as $N\to\infty$ with fixed large $T$. Its proof also implies that $\widehat{s}/T$ consistently estimates $s_{0}/T$ as both $N,T\to\infty$.
(ii) When the columns of $\mathbf{F}$} have distinct sparsity parameters, it is an important step to estimate the largest one among $\{s_{1},...,s_{r}\}$ first, and $\widehat{s}$ in Theorem 4 is an estimator for $s^{*}=\max\{s_{1},...,s_{r}\}$. In other words, for large $T$, we can show that
$\lim_{N\rightarrow\infty}P(\widehat{s}=s^{*})=1.$ (23)
@@
(ii) In practice, there are many choices for $g(N_{1},T)$ in (14). For example, we may take $C_{T}=\log(T)$, $g(N_{1},T)=\frac{\sqrt{N_{1}}+T}{\sqrt{N_{1}}T}\log(\frac{\sqrt{N_{1}}T}{\sqrt{N_{1}}+T})$. By an argument similar to that in Corollary 1 of Bai and Ng (2002),
$IC(s)=\ln(R^{J}(s))+r\frac{s}{T}\frac{\sqrt{N_{1}}+T}{\sqrt{N_{1}}T}\log(T)\log(\frac{\sqrt{N_{1}}T}{\sqrt{N_{1}}+T})$ (24)
can also consistently estimate $s_{0}$, where $r$ can be replaced by $\widehat{r}$ obtained in (12).